# Position-building in competition with real-world constraints

### v1

@@ 6.2 Plots of unconstrained two-trader equilibrium
In the following plots we examine two-trader equilibria for various values of $\kappa$ and $\lambda$ with $\gamma=0.2$ in all cases. At this stage we have no investigated what the optimal value for $\gamma$ is but illustrate in the sequel that the algorithm proposed has instances that converge with $\gamma<1$ but not with $gamma=1$.
@@
Figure 9: Two-trader equilibria solutions compared with the closed for solutions in (4.6a) and (4.6b). The solutions presented are the reconstructed functions $a^{*}$ and $b^{*}$ in Eq. 5.9 that arise out of the iterative quadratic program Algorithm 1 with $\gamma=0.2$ then compared with the closed-form solutions $a_{\texttt{eq}}$ and $b_{\texttt{eq}}$ that arise from the analytical solutions in [Chriss(2024)] and which are presented in equations (4.6a) and (4.6b). In this plot we solve for the two-trader equilibria for $\kappa=1$ and $\lambda=5$ using 20 Fourier terms. The upper left-hand plot shows $a^{*}$ traded along with $a_{\texttt{eq}}$ and $b^{*}$ along with $b_{\texttt{eq}}$, as well as the $L_{2}$ norms of the differences between the theoretical equilibrium strategies and their approximations. The upper right-hand shows the difference between the reconstructed functions and theoretical equilibrium functions at the first and last iterations of the optimization. The state-space diagram in the lower-left plot shows how the total cost of trading for $a$ and $b_{\lambda}$ plotted as coordinates starting with the total cost of trading directly prior to Step 1(i) of the algorithm which in this case means both traders trade a risk-neutral strategy. We see that the equilibrium strategies substantially benefit $a$ by decreasing the total cost of trading from approximately 9 to approximately 8.2, while increasing the costs of $b_{\lambda}$ modestly from 45 to approximately 46.2.
@@
Figure 10: Two-trader equilibria solutions compared with the closed-form for $\lambda=20,\kappa=6$ and $N=25$ Fourier series terms for Algorithm 1 with $\gamma=0.2$. The solutions presented are the reconstructed functions $a^{*}$ and $b^{*}$ in Eq. 5.9 and are compared with the closed-form solutions $a_{\texttt{eq}}$ and $b_{\texttt{eq}}$ in equations (4.6a) and (4.6b). The solver converges after 99 iterations, and the top left chart shows the $L_{2}$ norms of the differences between the theoretical equilibrium strategies and their approximations.. The state-space diagram plots the cost of trading for $a$ and $b_{\lambda}$ starting with the initial point (labeled init) which represents both traders trading the risk-neutral strategy. In this case the $b_{\lambda}$ trades 20 times as much as $a$ and $a$’s initial best-response to $b_{\lambda}=\lambda\cdot t$ already substantially reduces its cost while increasing $b_{\lambda}$’s.
@@
Figure 11: Unconstrained two-trader equilibria solutions compared with the closed-form for $\lambda=1,\kappa=25$ and $N=35$ terms in the Fourier series approximations for Algorithm 1 with $\gamma=0.2$. The solutions presented are the reconstructed functions $a^{*}$ and $b^{*}$ in Eq. 5.9 and are compared with the closed-form solutions $a_{\texttt{eq}}$ and $b_{\texttt{eq}}$ in equations (4.6a) and (4.6b). The solver converges after 62 iterations with $\gamma=0.2$, and the top left chart shows the $L_{2}$ norms of the differences between the theoretical equilibrium strategies and their approximations. The state-space diagram plots the cost of trading for $a$ and $b_{\lambda}$ starting with the initial point (labeled init) which represents both traders trading the risk-neutral strategy. Of interest is that the total cost of trading is higher for both traders at equilibrium than it would be if they both traded a risk-neutral strategy as they do at the point labeled init guess. That is, they would both be better off if they could coordinate.
Ensemble plots of unconstrained two-trader equilibria: In these plots we show various plots of interest along with state-space diagrams. In all of these plots $\gamma=0.2$ (see Algorithm 1).
@@
Figure 12: Unconstrained two-trader equilibria solutions compared with the closed-form for a cross-section of parameters using Algorithm 1 with $\gamma=0.2$. The solutions presented are the reconstructed functions $a^{*}$ and $b^{*}$ in Eq. 5.9 and are compared with the closed-form solutions $a_{\texttt{eq}}$ and $b_{\texttt{eq}}$ in equations (4.6a) and (4.6b). The number of terms $N$ for each set of $\lambda$ and $\kappa$ is chosen to ensure a given level of approximation accuracy, which we measure by the $L_{2}$ norm of the difference between the exact analytic and the approximated trading strategy (shown above each top chart). The state-space diagram in the bottom row plots the cost of trading for $a$ and $b_{\lambda}$ starting with the initial point (labelled init guess), which represents both traders trading the risk-neutral strategy. Of interest is for the two rightmost charts the total cost of trading is higher for both traders than it would be if they both traded a risk-neutral strategy as in initiation – they would both be better off, but because collusion is prohibited, they inevitably both end up trading the equilibrium strategies, which are more expensive than trading the risk-neutral strategies.
@@
Figure 13: Unconstrained two-trader equilibria solutions compared with the closed-form for a cross-section of parameters using Algorithm 1 with $\gamma=0.2$. The sets of parameters $\lambda$, $\kappa$ is the same as in the previous chart 12, but the number of Fourier terms is fixed at lower $N=10$ for all charts. Both the $L_{2}$ norms of the approximation error and visual inspection illustrate the dependence of the approximation accuracy on the number of Fourier terms $n$. For example, for $\kappa=20$, $\lambda=1$ parameters (second chart from the right) the $L_{2}$ approximation error for $N=10$ is almost 12 times higher than for $n=30$ (0.0025 vs. 0.0002).
Figure 14: Unconstrained two-trader equilibria solutions compared with the closed-form for using Algorithm 1 with $\gamma=0.2$ and an increasing number of Fourier coefficients.
The impact of $\gamma$ on algorithm convergence: As explained in Algorithm 1 it is sometimes helpful to not move toward equilibrium strategies at each step of the algorithm but move along the lines that interpolates between the prior step’s strategy and the best-response to that. We have not explored how to optimally set $\gamma$ but in Figure 16 we demonstrate for the case $\kappa=25$ and $\lambda=1$ the algorithm diverges when $\gamma=1$ and converges with $\gamma\leq 0.6$.
Figure 15: A visualization of a single scenario for unconstrained two-0trader equilibrium for the case $\kappa=5$, $\lambda=1$ with a variety of values for $\gamma$ to illustrate the impact of the number of steps taken on the equilibrium path.
@@ Appendix A Proofs of cost function approximations
which leads to Eq. A.16d becoming
$\frac{\kappa}{2}-\kappa\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2a_{n}}{n\pi}+\kappa\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2a_{n}}{n\pi}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}nm\cdot\frac{a_{n}a_{m}}{m^{2}-n^{2}}$ $=\frac{\kappa}{2}+2\kappa\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}a_{m}n\frac{m}{m^{2}-n^{2}}$ (A.20)
@@
$\int_{0}^{1}\kappa\,\dot{a}_{\mathcal{F}}a_{\mathcal{F}}=\frac{\kappa}{2}+2\kappa\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}a_{m}n\frac{m}{m^{2}-n^{2}}$ (A.21)
@@
$2\kappa\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\!\!a_{n}a_{m}n\frac{m}{m^{2}-n^{2}}+2\kappa\lambda\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}b_{m}n\cdot\frac{m}{m^{2}-n^{2}}$ $=2\kappa\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\left\{\frac{a_{n}a_{m}nm}{m^{2}-n^{2}}+\frac{\lambda a_{n}b_{m}nm}{m^{2}-n^{2}}\right\}$ (A.35)
$=2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\left(a_{m}+\lambda b_{m}\right)\frac{a_{n}\,n\,m}{m^{2}-n^{2}}$ (A.36)
@@
$\begin{split}&\text{Cost}_{\mathcal{F},a}(b_{\lambda})=\\ &\quad\frac{(2+\kappa)(1+\lambda b_{\lambda})}{2}+\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}(a_{n}^{2}+\lambda a_{n}b_{n})+\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(b_{n}-a_{n})}{n\pi}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\left(a_{m}+\lambda b_{m}\right)\frac{a_{n}\,n\,m}{m^{2}-n^{2}}\end{split}$ (A.37)
@@
$\begin{split}\text{Cost}_{\mathcal{F},b_{\lambda}}(a)&=\lambda+\lambda^{2}+\frac{\kappa\lambda}{2}+\frac{\kappa\lambda^{2}}{2}+\\ &\quad\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}\left(a_{n}b_{n}+\lambda b_{n}^{2}\right)+\\ &\quad\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(a_{n}-b_{n})}{n\pi}+\\ &\quad 2\kappa\lambda\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}a_{m}n\cdot\frac{m}{m^{2}-n^{2}}+2\kappa\lambda^{2}\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}b_{m}n\frac{m}{m^{2}-n^{2}}\end{split}$ (A.48)
@@
$\begin{split}\text{Cost}_{\mathcal{F},b_{\lambda}}(a)&=\lambda+\lambda^{2}+\frac{\kappa\lambda}{2}+\frac{\kappa\lambda^{2}}{2}+\\ &\quad\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}\left(a_{n}b_{n}+\lambda b_{n}^{2}\right)+\\ &\quad 2\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{a_{n}-b_{n}}{n\pi}+\\ &\quad 2\kappa\lambda\left(\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}a_{m}n\cdot\frac{m}{m^{2}-n^{2}}+\lambda\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}b_{m}n\frac{m}{m^{2}-n^{2}}\right)\end{split}$ (A.49)
finally noting that for any pair of $m,n$ such that $n+m$ is odd we have
$b_{n}a_{m}n\frac{m}{m^{2}-n^{2}}+\lambda b_{n}b_{m}n\frac{m}{m^{2}-n^{2}}=b_{n}n\frac{m}{m^{2}-n^{2}}(a_{m}+\lambda b_{m})$ (A.50)
the last expression in Eq. A.49 becomes
$\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}a_{m}n\cdot\frac{m}{m^{2}-n^{2}}+\lambda\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}b_{m}n\frac{m}{m^{2}-n^{2}}=\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{b_{n}nm}{m^{2}-n^{2}}(a_{m}+\lambda b_{m})$ (A.51)
@@
$\begin{split}\lambda+\lambda^{2}+\frac{\kappa\lambda}{2}&=\lambda(1+\lambda++\frac{\kappa}{2}+\frac{\kappa\lambda}{2})\\ &=\lambda\frac{2(1+\lambda b_{\lambda})+\kappa(1+\lambda b_{\lambda})}{2}\\ &=\frac{\lambda(1+\lambda b_{\lambda})(2+\kappa)}{2}\end{split}$ (A.52)
@@
$\text{Cost}_{\mathcal{F},b_{\lambda}}(a)=\frac{\lambda(1+\lambda b_{\lambda})(2+\kappa)}{2}+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}\left(a_{n}b_{n}+\lambda b_{n}^{2}\right)+2\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{(a_{n}-b_{n})}{n\pi}+2\kappa\lambda\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{b_{n}nm}{m^{2}-n^{2}}(a_{m}+\lambda b_{m})$ (A.53)
The conclusion after some re-arrangement to make the formula conform to $\text{Cost}_{\mathcal{F},a}(b_{\lambda})$ we obtain:
$\begin{split}&\text{Cost}_{\mathcal{F},b_{\lambda}}(a)=\\ &\quad\lambda\left(\frac{1}{2}(2+\kappa)(1+\lambda b_{\lambda})+\frac{\pi^{2}}{2}\sum_{n=1}^{N}\left(\lambda b_{n}^{2}+a_{n}b_{n}\right)n^{2}+\frac{2\kappa}{\pi}\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{a_{n}-b_{n}}{n}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{(a_{m}+\lambda b_{m})b_{n}nm}{m^{2}-n^{2}}\right)\end{split}$ (A.54)

### v2

@@ 6.2 Plots of unconstrained two-trader equilibrium
In the following plots we examine two-trader equilibria for various values of $\kappa$ and $\lambda$ with $\gamma=0.2$ in all cases. At this stage we have no investigated what the optimal value for $\gamma$ is but illustrate in the sequel that the algorithm proposed has instances that converge with $\gamma<1$ but not with $\gamma=1$.
@@
Figure 9: Two-trader equilibria strategies $a^{*},b_{\lambda}^{*}$ compared with the closed-form solutions $a_{\texttt{eq}},b_{\texttt{eq}}$ in (4.6a) and (4.6b) for $\kappa=1$ and $\lambda=5$ using $N=20$ Fourier terms. The upper left-hand plot shows $a^{*}$ traded along with $a_{\texttt{eq}}$ and $b^{*}$ along with $b_{\texttt{eq}}$, as well as the $L_{2}$ norms of the differences between the theoretical equilibrium strategies and their approximations. The upper right-hand shows the difference between the reconstructed functions and theoretical equilibrium functions at the first and last iterations of the optimization. The state-space diagram in the lower-left plot shows the total cost of trading for $a$ and $b_{\lambda}$ directly prior to Step 1(i) when both traders trade a risk-neutral strategy. The plot shows that the solution (the equilibrium strategy) benefits $a$, lowering the total cost trading from approximately 9 to 8.2, while increasing $b_{\lambda}$’s modestly 45 to approximately 46.2. The solver converged in 12 iterations and $\gamma=0.8$.
@@
Figure 10: Two-trader equilibria solutions compared with the closed-form for $\lambda=20,\kappa=6$ and $N=25$ Fourier series terms for Algorithm 1 with $\gamma=0.8$ and Algorithm 1 converged in 21 iterations. The solutions presented are the reconstructed functions $a^{*}$ and $b^{*}$ in Eq. 5.9 and are compared with the closed-form solutions $a_{\texttt{eq}}$ and $b_{\texttt{eq}}$ in equations (4.6a) and (4.6b). The top left chart shows the $L_{2}$ norms of the differences between the theoretical equilibrium strategies and their approximations.. The state-space diagram plots the cost of trading for $a$ and $b_{\lambda}$ starting with the initial point (labeled init) which represents both traders trading the risk-neutral strategy. In this case the $b_{\lambda}$ trades 20 times as much as $a$ and $a$’s initial best-response to $b_{\lambda}=\lambda\cdot t$ already substantially reduces its cost while increasing $b_{\lambda}$’s.
@@
Figure 11: Unconstrained two-trader equilibria solutions compared with the closed-form for $\lambda=1,\kappa=25$ and $N=35$ terms in the Fourier series approximations for Algorithm 1 with $\gamma=0.2$ and the algorithm converged in 62 iterations. The solutions presented are the reconstructed functions $a^{*}$ and $b^{*}$ in Eq. 5.9 and are compared with the closed-form solutions $a_{\texttt{eq}}$ and $b_{\texttt{eq}}$ in equations (4.6a) and (4.6b). The top left chart shows the $L_{2}$ norms of the differences between the theoretical equilibrium strategies and their approximations. The state-space diagram plots the cost of trading for $a$ and $b_{\lambda}$ starting with the initial point (labeled init) which represents both traders trading the risk-neutral strategy. Of interest is that the total cost of trading is higher for both traders at equilibrium than it would be if they both traded a risk-neutral strategy as they do at the point labeled init guess. That is, they would both be better off if they could coordinate. It’s also interesting that the convergence is much slower for higher $\gamma$, for example if we use the same value of $\gamma=0.8$ as in the previous example, the solver does converge within tolerance even after 100 iterations.
Ensemble plots: In Figures 13 to 17 we depict "ensembles" of plots of two-trader equilibrium to demonstrate how a single parameter’s change influences the optimization results. For example, as explained in Algorithm 1 it can be helpful to advance in each step using convex combinations of prior- and next-step strategies. We have not explored how to optimally set $\gamma$ but in Figure 16 we demonstrate for the case $\kappa=25$ and $\lambda=1$ the algorithm diverges when $\gamma=1$ and converges when $\gamma\leq 0.6$.
@@
Figure 12: Unconstrained two-trader equilibria solutions compared with the closed-form for a cross-section of parameters using Algorithm 1 with the values of $\gamma$ selected to ensure quick convergence (i.e. $\gamma=0.8$ for the left two plots, and $\gamma=0.2$ for the right three). The solutions presented are the reconstructed functions $a^{*}$ and $b^{*}$ in Eq. 5.9 and are compared with the closed-form solutions $a_{\texttt{eq}}$ and $b_{\texttt{eq}}$ in equations (4.6a) and (4.6b). The number of terms $N$ for each set of $\lambda$ and $\kappa$ is chosen to ensure a given level of approximation accuracy, which we measure by the $L_{2}$ norm of the difference between the exact analytic and the approximated trading strategy (shown above each top chart). The state-space diagram in the bottom row plots the cost of trading for $a$ and $b_{\lambda}$ starting with the initial point (labelled init guess), which represents both traders trading the risk-neutral strategy. Of interest is for the two rightmost charts the total cost of trading is higher for both traders than it would be if they both traded a risk-neutral strategy as in initiation – they would both be better off, but because collusion is prohibited, they inevitably both end up trading the equilibrium strategies, which are more expensive than trading the risk-neutral strategies.
@@
Figure 13: Unconstrained two-trader equilibria solutions compared with the closed-form for a cross-section of parameters using Algorithm 1 with the values of $\gamma$ selected to ensure quick convergence (i.e. $\gamma=0.8$ for the left two plots, and $\gamma=0.2$ for the right three). The sets of parameters $\lambda$, $\kappa$, $\gamma$ is the same as in the previous chart 12, but the number of Fourier terms is fixed at lower $N=10$ for all charts. Both the $L_{2}$ norms of the approximation error and visual inspection illustrate the dependence of the approximation accuracy on the number of Fourier terms $N$. For example, for $\kappa=20$, $\lambda=1$ parameters (second chart from the right) the $L_{2}$ approximation error for $N=10$ is almost 12 times higher than for $N=30$ (0.0025 vs. 0.0002).
Figure 14: This plot demonstrates the impact of a the number of Fourier coefficients $N$ used in forming the approximate cost functions (see Propositions 2.1 and 2.2) for the case $\kappa=20,\lambda=5$ and $\gamma=0.5$.
Figure 15: Unconstrained two-trader equilibrium for $\kappa=5$, $\lambda=1$ for $\gamma$ increasing from left to right from 0.2 to 1.0; higher $\gamma$ decreases the number of iterations required for convergence, while higher $\gamma$ eventually causes the solver to diverge.
@@ Appendix A Proofs of cost function approximations
and we note that this term is zero because for every pair $(m,n),n+m\text{odd}$, the is a corresponding pair $(m,n)$ which cancels. This leads to Eq. A.16d becoming
$\frac{\kappa}{2}-\kappa\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2a_{n}}{n\pi}+\kappa\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2a_{n}}{n\pi}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}nm\cdot\frac{a_{n}a_{m}}{m^{2}-n^{2}}$ $=\frac{\kappa}{2}$ (A.20)
@@
$\int_{0}^{1}\kappa\,\dot{a}_{\mathcal{F}}a_{\mathcal{F}}=\frac{\kappa}{2}$ (A.21)
@@
$2\kappa\lambda\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}a_{n}b_{m}n\cdot\frac{m}{m^{2}-n^{2}}$ (A.35)
@@
$\text{Cost}_{\mathcal{F},a}(b_{\lambda})=\frac{(2+\kappa)(1+\lambda)}{2}+\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}(a_{n}^{2}+\lambda a_{n}b_{n})+\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(b_{n}-a_{n})}{n\pi}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{\lambda b_{m}a_{n}\,n\,m}{m^{2}-n^{2}}$ (A.36)
@@
$\begin{split}\text{Cost}_{\mathcal{F},b_{\lambda}}(a)&=\lambda+\lambda^{2}+\frac{\kappa\lambda}{2}+\frac{\kappa\lambda^{2}}{2}+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}\left(a_{n}b_{n}+\lambda b_{n}^{2}\right)+\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{2(a_{n}-b_{n})}{n\pi}+\\ &\quad 2\kappa\lambda\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}a_{m}n\cdot\frac{m}{m^{2}-n^{2}}+2\kappa\lambda^{2}\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}b_{m}n\frac{m}{m^{2}-n^{2}}\end{split}$ (A.47)
@@
$\begin{split}\text{Cost}_{\mathcal{F},b_{\lambda}}(a)&=\lambda+\lambda^{2}+\frac{\kappa\lambda}{2}+\frac{\kappa\lambda^{2}}{2}+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}\left(a_{n}b_{n}+\lambda b_{n}^{2}\right)+2\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{a_{n}-b_{n}}{n\pi}+\\ &\quad 2\kappa\lambda\left(\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}a_{m}n\cdot\frac{m}{m^{2}-n^{2}}+\lambda\!\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}b_{m}n\frac{m}{m^{2}-n^{2}}\right)\end{split}$ (A.48)
finally noting that the last is zero due to cancellations, we have: the last expression in Eq. A.48 becomes
$2\kappa\lambda\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}b_{n}a_{m}n\cdot\frac{m}{m^{2}-n^{2}}$ (A.49)
@@
$\begin{split}\lambda+\lambda^{2}+\frac{\kappa\lambda}{2}&=\lambda\left(1+\lambda++\frac{\kappa}{2}+\frac{\kappa\lambda}{2}\right)\\ &=\lambda\frac{2(1+\lambda)+\kappa(1+\lambda)}{2}\\ &=\frac{\lambda(1+\lambda)(2+\kappa)}{2}\end{split}$ (A.50)
@@
$\text{Cost}_{\mathcal{F},b_{\lambda}}(a)=\frac{\lambda(1+\lambda)(2+\kappa)}{2}+\lambda\frac{\pi^{2}}{2}\sum_{n=1}^{N}n^{2}\left(a_{n}b_{n}+\lambda b_{n}^{2}\right)+2\kappa\lambda\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{(a_{n}-b_{n})}{n\pi}+2\kappa\lambda\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{a_{m}b_{n}nm}{m^{2}-n^{2}}$ (A.51)
The conclusion after some re-arrangement and noting that the last term with $\lambda b_{n}b_{m}nm$ cancel so that we arrive at the formula conform to $\text{Cost}_{\mathcal{F},a}(b_{\lambda})$ in Proposition 2.2:
$\text{Cost}_{\mathcal{F},b_{\lambda}}(a)=\lambda\left\{\frac{1}{2}(2+\kappa)(1+\lambda)\!+\!\frac{\pi^{2}}{2}\sum_{n=1}^{N}\left(\lambda b_{n}^{2}+a_{n}b_{n}\right)n^{2}\!+\!\frac{2\kappa}{\pi}\sum_{\begin{subarray}{c}n=1\\ n\,\text{odd}\end{subarray}}^{N}\frac{a_{n}-b_{n}}{n}+2\kappa\!\!\!\!\sum_{\begin{subarray}{c}n,m=1\\ n+m\,\text{odd}\end{subarray}}^{N}\frac{a_{m}b_{n}nm}{m^{2}-n^{2}}\right\}$ (A.52)